OpenStudy (anonymous):

Can you please explain how to find Domain, Range, Increase, Decrease. For example y=5/x+3

OpenStudy (anonymous):

5/x + 3 is just a vertical translation of 5/x 3 units up... so their domain is the same. their range is different... 5/x range is (-inf, 0) U (0, inf). so the range for 5/x+3 is (-inf,3) U (3,inf). increase and decrease intervals are the same for 5/x and 5/x+3...

OpenStudy (anonymous):

i'm assuming this is the function y = (5/x) + 3

OpenStudy (anonymous):

i think the range is (-inf,0) and (0,nif)

OpenStudy (anonymous):

that all depends... is your equation y = 5/(x+3) or y = 5/x + 3

OpenStudy (anonymous):

oh ok. mine is y=5/(x+3)

OpenStudy (anonymous):

if it's the second one, the range is (-inf, 3) U (3, inf)

OpenStudy (anonymous):

then forget everything i said....

OpenStudy (anonymous):

thank you) what is domain in both cases?